OpenStudy (anonymous):

9y/y-10 + 1 = 8y/y-10

OpenStudy (anonymous):

y=?

OpenStudy (anonymous):

y=0 or y=10.

OpenStudy (lalaly):

\[\frac{9y}{y-10}-\frac{8y}{y-10}=-1\]\[\frac{y}{y-10}=-1\]\[y=-y+10\]\[2y=10\]\[y=5\]

OpenStudy (anonymous):

9y/y-10 + 1 = 8y/y-10 9y/y-10 -8y/y-10=-1 (9y-8y)/y-10=-1 y/y-10=-1 y=10-y 2y=10 y=5